OpenStudy (anonymous):

Very hard indefinite integral cosx*(sinx)^(1/2)

OpenStudy (turingtest):

try u=sinx

OpenStudy (anonymous):

So u = cos x and dv = sin^0.5 du = -sin x and v, here is my uncertainty is v = -cos^(1/2)? if so then u'v+uv' = -sinx*sinx^1/2 +cosx *-cosx ^1/2

OpenStudy (anonymous):

where do u get sin x from?

OpenStudy (turingtest):

by parts not necessary, or even possible here simple u sub I got the sinx from inside the (sinx)^1/2

OpenStudy (anonymous):

\[\int\limits_{}^{} \cos x * \sqrt{sinx}\]

OpenStudy (turingtest):

indeed, let u=sinx then du=....

OpenStudy (anonymous):

okey du is then cos x

OpenStudy (turingtest):

so the integral is then what? (in terms of u)

OpenStudy (anonymous):

integral of u*dv ?

OpenStudy (turingtest):

why do you insist on trying to integrate by parts? there is no dv

OpenStudy (anonymous):

or u'v+uv' = (uv)'

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

sin x * -cos x?

OpenStudy (turingtest):

\[\int\sqrt{\sin x}\cos xdx\\u=\sin x\\du=\cos xdx\]rewrite in terms of \(u\)

OpenStudy (turingtest):